3. Two capacitors Ci and C are connected in parallel. The total charge of the capacitors is O. What is the voltage on the system? What are the charges of the capacitors? Calculate numerical values for V-6 V and C
Answer #1

equivalent capacitance will be:

Ceq = C1 + C2

Net charge = Q

We know that:

Q = Ceq*V

V = Q/Ceq

V = Q/(C1 + C2)

Now in parallel circuit voltage on each capacitor will be same, So

V1 = V2 = V

Q1 = C1*V1 = C1*V

Q1 = C1*Q/(C1 + C2)

Similarly

Q2 = C2*Q/(C1 + C2)

for given values:

Q = Ceq*V

Q = (1 + 3)*6 = 24 uC

Q1 = C1*V = 1*6 = 6 uC

Q2 = C2*V = 3*6 = 18 uC